Find class boundaries, midpoint, and width for the class 122-128

1
Expert's answer
2022-03-20T06:43:38-0400

Class boundaries: 121.5 - 128.5.

Midpoint: "\frac{122+128}{2}=125."

Width: 128.5-121.5=7.